# Leveraging Large Language Models to Detect Missed Peephole Optimizations

### v1

@@ 12        $\textnormal{{dedup\_set}}\leftarrow\textnormal{{dedup\_set}}\cup\{\textnormal{{digest}}\}$
13        $\textnormal{{result}}\leftarrow\textnormal{{result}}\cup\{\textnormal{{seq}}\}$
@@ 25    $\textnormal{{new\_set}}\leftarrow\varnothing$
26     foreach instruction sequence $\textnormal{{seq}}\in\textnormal{{seq\_set}}$ do
27        if any instruction in seq depends on inst then
28           $\textnormal{{new\_set}}\leftarrow\textnormal{{new\_set}}\cup\{[\textnormal{{inst}}]+\textnormal{{seq}}\}$
@@ 29           $\textnormal{{added}}\leftarrow\textnormal{{true}}$
30        else  $\textnormal{{new\_set}}\leftarrow\textnormal{{new\_set}}\cup\{\textnormal{{seq}}\}$
@@ 33     $\textnormal{{seq\_set}}\leftarrow\textnormal{{new\_set}}$
Algorithm 1 presents the design of Lampo’s extractor, whose goal is to obtain all unique dependent instruction sequences from each basic block in every LLVM IR module in the corpus. The extractor takes two inputs: an LLVM IR module and a set, dedup_set containing all previously seen encoded instruction sequences for deduplication. It outputs a set of unique instruction sequences extracted from the given module. Each of the instruction sequences is wrapped as an LLVM function and passed to the LLM to explore potential optimization opportunity within it. The detailed steps are described below.
@@
Extracting From Basic Blocks.  Given an LLVM IR module, Lampo first initializes result to store the extracted sequences (Algorithm 1). It then traverses each basic block bb of each function in the module (Algorithms 11), and extracts all dependent instruction sequences contained within bb by calling function ExtractSeqsFromBB (Algorithm 1). Specifically, ExtractSeqsFromBB traverses the instructions in bb in reverse order (Algorithm 1). For each instruction inst, if it is a terminator (i.e., the last instruction in a basic block), then it is skipped (Algorithm 1), since terminators only define control flow jumps, and Lampo currently targets only missed optimization within basic blocks. On Algorithm 1, Lampo initializes a flag added as false to indicate whether inst has been inserted into any sequence, and initialize a new set, new_set, which is eventually used to update seq (Algorithm 1). Next, for each instruction sequence seq in seq_set, if any instruction in seq depends on inst, Lampo prepends inst to seq and sets added to true (Algorithms 11). Otherwise, seq remains unchanged (Algorithm 1). If inst is not added to any existing sequences—meaning no traversed instruction depends on instLampo creates a new sequence containing only inst and adds the new sequence to new_set (Algorithms 11).
@@
Verifying and Deduplicating Instruction Sequences.  After Lampo traverses all the instructions within a basic block, a set of instruction sequences can be obtained. For each extracted instruction sequence, Lampo wraps it as an LLVM IR function (Algorithm 1). Specifically, Lampo appends a return instruction that returns the value produced by the last instruction to the sequence and treats all undefined operands in the sequence of instructions as function arguments. With the wrapped function obtained, Lampo checks that LLVM cannot further optimize the instruction sequence (Algorithm 1). This check is necessary because a sequence that is unoptimizable in its original context may become optimizable once isolated. For example, an instruction in the sequence might be used not only by subsequent instructions within the sequence but also by instructions in later basic blocks; once those external uses are removed, the instruction could be combined with others in the sequence. Finally, Lampo computes a hash of the function based on the opcode and operands of each instruction (Algorithm 1), and add instruction sequences whose hashes have not been recorded in dedup_set (Algorithms 11).
@@ 3.3. Verifying Outputs of the LLM
After extracting instruction sequences from the corpus and wrapping them as LLVM IR functions (we use the term instruction sequence and function interchangeably in the following), Lampo begins prompting the specified LLM to optimize each function. Each LLM-generated candidate is then verified through three steps.
@@
Preprocessing with opt.  First, Lampo feeds the candidate to opt, the LLVM optimizer, with the -O3 flag to perform aggressive optimization (step 3). The step serves two purposes. First, opt checks whether the function proposed by the LLM is syntactically correct. If the function contains syntax errors, opt outputs an error message, which Lampo can then use as feedback to prompt the LLM to fix the error (step 6). Second, although the LLM may produce a modified version of the original function, the proposed function may not be canonical or optimal w.r.t. opt. In such cases, running opt can further optimize and canonicalize the proposed function reducing the effort required for subsequent manual analysis.
@@
Checking Interestingness.  If the function proposed by the LLM is syntactically correct, Lampo next checks whether the function—after preprocessed by opt — potentially manifests a beneficial optimization. Determining whether the transformation from the original function to the proposed one is beneficial is tricky, because (1) comparison between two functions is not straightforward, as multiple aspects can be considered (e.g., code size and performance on a specific target); (2) even if the proposed function appears significantly better, implementing the corresponding optimization might not be beneficial overall if it hinders other, more valuable optimizations. Due to these challenges, this step is designed to only filter out as many uninteresting cases as possible (i.e., those unlikely to manifest a beneficial optimization), thereby reducing the effort required for subsequent manual analysis. In our prototype, two metrics are considered when checking the interestingness, the instruction count and the total cycles measured by llvm-mca  [^1] A performance analysis tool to statically measures the performance of machine code in a specific CPU based on LLVM’s scheduling models. on a specific target with a specific CPU (e.g., with the target triple set to x86_64-unknown-unknown and the CPU set to btver2) Functions with fewer instructions or fewer total cycles are considered interesting and thus pass the check. Additionally, functions with the same instruction count and total cycles are also considered interesting if they differ syntactically from the original, since such transformations—though not immediately improving the function—may enable further optimization. Interestingness checking is performed before correctness verification, as it is generally more efficient, avoiding the costlier correctness checking for unpromising cases and thus improving overall workflow efficiency.
@@
Verifying Correctness.  If the function proposed by the LLM passes the interestingness checking, Lampo then verifies the correctness of the transformation from the original function to the proposed one using Alive2 (Lopes et al., 2021). If Alive2 proves the transformation incorrect, it provides a counterexample to demonstrate the incorrectness, which Lampo uses as feedback to prompt the LLM to propose a new candidate. If Alive2 encounters a fixable error that prevents the proof—for example, mismatched function signatures—the error message is also used as feedback to guide the LLM in avoiding the issue (step 6). Otherwise, if Alive2 proves the transformation correct, Lampo records the pair of functions as a potential missed optimization (step 7)
@@ 4. Evaluation
To comprehensively evaluate Lampo, we conducted experiments to investigate the following research questions.
RQ1: Can Lampo detect previously reported missed optimizations?
RQ2: Can Lampo discover new missed optimizations?
RQ3: How much throughput can Lampo achieve, and what is the associated cost?
@@ 4.1. Experimental Setup
Selected LLM Models.Table 1 lists all the models selected for evaluating Lampo. We selected diverse models to investigate how the capabilities of different LLMs affects the effectiveness of Lampo. For RQ1, we selected two open-source and four proprietary models. Specifically, we chose two open-source models with significantly different sizes: Gemma3 (27 billion parameters) and Llama3.3 (70 billion parameters). For proprietary models, we included two base models, i.e., Gemini2.0 and GPT-4.1, and two reasoning models, i.e., Gemini2.0T and o4-mini. These models have relatively early knowledge cut-off dates, allowing us to collect a reasonable number of benchmarks while avoiding potential data leakage. In RQ2, during the long-term, intermittent experiment, we tried different open-source models, and a locally deployed version of llama3.3:70b is the most used one. In RQ3, Llama3.3 and Gemini2.5 (a cost-efficient and low-latency model recommended for scenarios that requires high throughput) are used to evaluate the throughput and cost, covering two common usage scenarios, i.e., using a locally deployed LLM and invoking an LLM via API. For Gemini2.0T and Gemini2.5, we set the reasoning budget to low (i.e., at most 1,024 reasoning tokens).
@@
Baselines.  We used two baselines in our evaluations.
@@
  • Lampo-. To evaluate the effectiveness of the iterative prompting strategy, we compared Lampo with a variant named Lampo-, which does not prompt the LLM to generate new candidates using feedback (i.e., it omits steps 4 and 7 in Figure 2)
@@ 4.2. RQ1: Detecting Previously Reported Missed Optimizations
Issue ID Gemma3 Llama3.3 Gemini2.0 Gemini2.0T GPT-4.1 o4-mini Souper
Lampo- Lampo Lampo- Lampo Lampo- Lampo Lampo- Lampo Lampo- Lampo Lampo- Lampo Default Enum
@@
Average 0.8 1.2 5.2 7 5.8 7.4 10.4 15.6 2.2 7.4 10.0 14.2 N/A N/A
Total 2 3 6 7 7 11 14 21 7 12 14 18 3 14
To quantitatively evaluate the effectiveness of Lampo in finding missed optimization opportunities and to objectively compare it with the baselines, we conducted a controlled experiment to assess how effectively each tool identifies previously reported missed optimizations. For the benchmark suite, we collected 25 missed peephole optimizations from issues in the official LLVM GitHub repository tagged with both missed-optimization and llvm:instcombine labels. To avoid potential data leakage (Zhou et al., 2025; López et al., 2025), we only included issues created after the end of August 2024, which is the latest knowledge cut-off date among all LLMs listed in Table 1 except for Gemini2.5, the model excluded from this research question. For each benchmark, we collected the example suboptimal LLVM IR function provided in the issue and examine whether Lampo, Lampo- and Souper could detect the missed optimization within the function. To mitigate the effects of the inherent non-determinism of LLMs, we repeated each experiment five times for both Lampo and Lampo-. The results of the experiment are shown in Table 2.
Effectiveness of Lampo with Different LLMs.  The effectiveness of Lampo heavily depends on the code understanding and reasoning capabilities of the employed LLM, and LLMs with different capabilities can therefore lead to varying levels of effectiveness. As shown in Table 2, the performance of Lampo varies significantly across different LLMs. With Gemma3, the smallest and weakest of the selected models, Lampo can detect at most 3 out of 25 missed optimizations, and on average only 1.2 per round. With other non-reasoning but more capable models, the performance of Lampo improves substantially: using Llama3.3, Gemini2.0, and GPT-4.1, Lampo can identify up to 7, 11 and 12 missed optimizations, and on average 7.0, 7.4, and 7.4, respectively. Employing reasoning models further enhance the performance of Lampo. With Gemini2.0T and o4-mini, Lampo can detect up to 21 and 18 missed optimizations, and on average 15.6 and 14.2, respectively.
@@
Lampo vs. Lampo-.  The iterative prompting strategy improves Lampo’s effectiveness in identifying missed optimizations regardless of the model employed. Without iterative prompting, Lampo- failed to identify 1 to 7 optimization that Lampo can detect across different LLMs. Moreover, the average number of optimizations detected by Lampo- is consistently lower than that of Lampo, with the gaps ranging from 0.4 to 5.2.
Table 3. All missed optimizations found by Lampo and reported to LLVM; the Can Souper detect? column indicates whether Souper can detect each of them. The result “timeout” means Souper cannot detect the missed optimization within a 20-minute timeout with Enum set to any value from 1 to 3.
Issue ID Status Can Souper detect?
Default Enum
@@ 4.3. RQ2: Discovering New Missed Optimization
To demonstrate the practical value of Lampo, we intermittently ran an experiment that employs Lampo to search for new missed optimization in LLVM. In this experiment, the corpus used by Lampo is from an LLVM IR dataset named LLVM Opt Benchmark (Zheng, 2023). This dataset contains optimized IR files from 239 real world projects written in C, C++, or Rust. Since the entire benchmark is too massive to serve as our corpus, we built our corpus by selecting five popular projects for each programming language, i.e., cpython, ffmpeg, linux, openssl, redis, node, protobuf, opencv, z3, pingora, ripgrep, typst, uv, zed. From this corpus, Lampo extracted over 800,000 unique instruction sequences, eliminating around 8.7 million duplicates.
Table 3 lists all missed optimizations found by Lampo and reported to LLVM, and Table 4 presents a summary of reported missed optimizations. Within seven months of development and intermittent testing, Lampo has found 26 potential missed peephole optimizations, with 15 confirmed and 6 already fixed in LLVM. It also detected a bug in Alive2.  [^2] https://anonymous.for.review Among all the reported missed optimizations, we marked 2 of them as wontfix (i.e., Issue 3 and Issue 4 ) according to the feedback from the maintainers. The reasons are: (1) the optimization is typically handled by the backend and (2) the optimization would prevent other valuable optimizations.
Table 4. Summary of Reported Missed Optimizations
Status Reported Confirm Fixed Duplicate Wontfix
Count 26 15 6 1 2
@@
Case Study 1 (Figures 4(a) and 4(d)).  The src function loads two consecutive 16-bit values from a given address %0 and combines them into a single 32-bit value using zero extension (zext), left shift (shl), and an or instruction. This function can be optimized to directly loading a 32-bit value from the given address. LLVM originally missed this optimization because it did not support merging consecutive loads with different sizes. Specifically, the original unoptimized function loads three consecutive values—two 8-bit values and one 16-bit value. LLVM successfully merges the two 8-bit loads but fails to further merge them with the subsequent 16-bit load. Souper cannot identify this case, as it does not support load and getelementptr instructions.
Case Study 2 (Figures 4(b) and 4(e)).  In the src function, the input %0 is first clamped to be at least 1 by llvm.umax.i8, then doubled with a 1-bit left shift, and finally clamped again to be at least 16. In the tgt function, the initial clamp is removed. This simplification is valid because the second clamp to 16 subsumes the effect of the first clamp, which makes it redundant. Souper cannot identify this missed optimization because it does not support the intrinsic group llvm.umax.*.
Case Study 3 (Figures 4(c) and 4(f)).  In the src function, the input %0 is first checked for being an ordinary (non-NaN) value; if it is NaN, it is replaced with 0.0, otherwise, %0 is kept. The result is then compared against 1.0. In the tgt function, the check and replacement are removed, and %0 is directly compared with 1.0. This simplification is valid because the instruction fcmp oeq (ordered equal) already returns false if %0 is NaN, so explicitly substituting NaNs with 0.0 is redundant. Souper cannot identify this missed optimization as it does not support floating point instructions.
@@
double %0, double 0.000000e+$0$
%4 = fcmp oeq double %3, 1.000000e+$0$
@@
Figure 4. Three examples of confirmed missed optimizations found by Lampo that Souper is unable to detect.
4.4. RQ3: Throughput and Cost of Lampo

### v2

@@ 12        $\textnormal{{dedup\_set}}\leftarrow\textnormal{{dedup\_set}}\cup\{\textnormal{{digest}}\}$
13        $\textnormal{{result}}\leftarrow\textnormal{{result}}\cup\{\textnormal{{instSeq}}\}$
@@ 25    $\textnormal{{new\_set}}\leftarrow\varnothing$
26     foreach instruction sequence $\textnormal{{instSeq}}\in\textnormal{{seq\_set}}$ do
27        if any instruction in instSeq depends on inst then
28           $\textnormal{{new\_set}}\leftarrow\textnormal{{new\_set}}\cup\{[\textnormal{{inst}}]+\textnormal{{instSeq}}\}$
@@ 29           $\textnormal{{added}}\leftarrow\textnormal{{true}}$
30        else $\textnormal{{new\_set}}\leftarrow\textnormal{{new\_set}}\cup\{\textnormal{{instSeq}}\}$
@@ 33     $\textnormal{{seq\_set}}\leftarrow\textnormal{{new\_set}}$
Algorithm 2 presents the design of LPO’s extractor, whose goal is to obtain all unique dependent instruction sequences from each basic block in every LLVM IR module in the corpus. The extractor takes two inputs: an LLVM IR module and a set, dedup_set containing all previously seen encoded instruction sequences for deduplication. It outputs a set of unique instruction sequences extracted from the given module. Each of the instruction sequences is wrapped as an LLVM function and passed to the LLM to explore potential optimization opportunity within it. The detailed steps are described below.
@@
Extracting From Basic Blocks.  Given an LLVM IR module, LPO first initializes result to store the extracted sequences (Algorithm 2). It then traverses each basic block bb of each function in the module (Algorithms 22), and extracts all dependent instruction sequences contained within bb by calling function ExtractSeqsFromBB (Algorithm 2). Specifically, ExtractSeqsFromBB traverses the instructions in bb in reverse order (Algorithm 2). For each instruction inst, if it is a terminator (i.e., the last instruction in a basic block), then it is skipped (Algorithm 2), since terminators only define control flow jumps, and LPO currently targets only missed optimization within basic blocks. On Algorithm 2, LPO initializes a flag added as false to indicate whether inst has been inserted into any sequence, and initialize a new set, new_set, which is eventually used to update instSeq (Algorithm 2). Next, for each instruction sequence instSeq in seq_set, if any instruction in instSeq depends on inst, LPO prepends inst to instSeq and sets added to true (Algorithms 22). Otherwise, instSeq remains unchanged (Algorithm 2). If inst is not added to any existing sequences—meaning no traversed instruction depends on instLPO creates a new sequence containing only inst and adds the new sequence to new_set (Algorithm 2).
@@
Verifying and Deduplicating Instruction Sequences.  After LPO traverses all the instructions within a basic block, a set of instruction sequences can be obtained. For each extracted instruction sequence, LPO wraps it as an LLVM IR function (Algorithm 2). Specifically, LPO appends a return instruction that returns the value produced by the last instruction to the sequence and treats all undefined operands in the sequence of instructions as function arguments. With the wrapped function obtained, LPO checks that LLVM cannot further optimize the instruction sequence (Algorithm 2). This check is necessary because a sequence that is unoptimizable in its original context may become optimizable once isolated. For example, an instruction in the sequence might be used not only by subsequent instructions within the sequence but also by instructions in later basic blocks; once those external uses are removed, the instruction could be combined with others in the sequence. Finally, LPO computes a hash of the function based on the opcode and operands of each instruction (Algorithm 2), and add instruction sequences whose hashes have not been recorded in dedup_set (Algorithms 22).
@@ 3.3. Verifying Outputs of the LLM
After extracting instruction sequences from the corpus and wrapping them as LLVM IR functions (we use the term instruction sequence and function interchangeably in the following), LPO begins prompting the specified LLM to optimize each function. Each LLM-generated candidate is then verified through three steps.
@@
Preprocessing with opt.  First, LPO feeds the candidate to opt, the LLVM optimizer, with the -O3 flag to perform aggressive optimization (step 3). The step serves two purposes. First, opt checks whether the function proposed by the LLM is syntactically correct. If the function contains syntax errors, opt outputs an error message, which LPO can then use as feedback to prompt the LLM to fix the error (step 6). Second, although the LLM may produce a modified version of the original function, the proposed function may not be canonical or optimal w.r.t. opt. In such cases, running opt can further optimize and canonicalize the proposed function reducing the effort required for subsequent manual analysis.
@@
Checking Interestingness.  If the function proposed by the LLM is syntactically correct, LPO next checks whether the function—after preprocessed by opt—potentially manifests a beneficial optimization. Determining whether the transformation from the original function to the proposed one is beneficial is tricky, because (1) comparison between two functions is not straightforward, as multiple aspects can be considered (e.g., code size and performance on a specific target); (2) even if the proposed function appears significantly better, implementing the corresponding optimization might not be beneficial overall if it hinders other, more valuable optimizations. Due to these challenges, this step is designed to only filter out as many uninteresting cases as possible (i.e., those unlikely to manifest a beneficial optimization), thereby reducing the effort required for subsequent manual analysis. In our prototype, two metrics are considered when checking the interestingness, the instruction count and the total cycles measured by llvm-mca  [^1] A performance analysis tool to statically measure the performance of machine code in a specific CPU based on LLVM’s scheduling models. on a specific target with a specific CPU (e.g., with the target triple set to x86_64-unknown-unknown and the CPU set to btver2) Functions with fewer instructions or fewer total cycles are considered interesting and thus pass the check. Additionally, functions with the same instruction count and total cycles are also considered interesting if they differ syntactically from the original, since such transformations—though not immediately improving the function—may enable further optimization. Interestingness checking is performed before correctness verification, as it is generally more efficient, avoiding the costlier correctness checking for unpromising cases and thus improving overall workflow efficiency.
@@
Verifying Correctness.  If the function proposed by the LLM passes the interestingness checking, LPO then verifies the correctness of the transformation from the original function to the proposed one using Alive2 (Lopes et al., 2021). If Alive2 proves the transformation incorrect, it provides a counterexample to demonstrate the incorrectness, which LPO uses as feedback to prompt the LLM to propose a new candidate. If Alive2 encounters a fixable error that prevents the proof—for example, mismatched function signatures—the error message is also used as feedback to guide the LLM in avoiding the issue (step 6). Otherwise, if Alive2 proves the transformation correct, LPO records the pair of functions as a potential missed optimization (step 7)
@@ 4. Evaluation
To comprehensively evaluate LPO, we conducted experiments to investigate the following research questions.
RQ1: Can LPO detect prior missed optimizations?
RQ2: Can LPO discover new missed optimizations?
RQ3: How much throughput can LPO achieve, and what is the associated cost?
@@ 4.1. Experimental Setup
Selected LLM Models.Table 1 lists all the models selected for evaluating LPO. We selected diverse models to investigate how the capabilities of different LLMs affect the effectiveness of LPO. For RQ1, we selected two open-source and four proprietary models. Specifically, we chose two open-source models with significantly different sizes: Gemma3 (27 billion parameters) and Llama3.3 (70 billion parameters). For proprietary models, we included two base models, i.e., Gemini2.0 and GPT-4.1, and two reasoning models, i.e., Gemini2.0T and o4-mini. These models have relatively early knowledge cut-off dates, allowing us to collect a reasonable number of benchmarks while avoiding potential data leakage. In RQ2, during the long-term, intermittent experiment, we tried several open-source models and used a locally deployed llama3.3:70b for most of the time. In RQ3, Llama3.3 and Gemini2.5 (a cost-efficient and low-latency model recommended for scenarios requiring high throughput) are used to evaluate the throughput and cost, covering two common usage scenarios, i.e., using a locally deployed LLM and invoking an LLM via API. For Gemini2.0T and Gemini2.5, we set the reasoning budget to low (i.e., at most 1,024 reasoning tokens).
@@
Baselines.  We used three baselines in our evaluations.
@@
  • Minotaur. We also compared LPO with Minotaur (Liu et al., 2024), a recent superoptimizer for LLVM that employs program synthesis techniques and focuses on optimizing integer and floating-point SIMD code. Default settings were used for Minotaur in our experiments.
  • LPO-. To evaluate the effectiveness of the iterative prompting strategy, we prepared LPO-, a variant of LPO that does not prompt the LLM to generate new candidates using feedback (i.e., it omits steps 4 and 7 in Figure 2)
@@ 4.2. RQ1: Detecting Previously Reported Missed Optimizations
Issue ID Gemma3 Llama3.3 Gemini2.0 Gemini2.0T GPT-4.1 o4-mini Souper Minotaur
LPO- LPO LPO- LPO LPO- LPO LPO- LPO LPO- LPO LPO- LPO Default Enum Default
@@
Average 0.8 1.2 5.2 7 5.8 7.4 10.4 15.6 2.2 7.4 10.0 14.2 N/A N/A N/A
Total 2 3 6 7 7 11 14 21 7 12 14 18 3 14 3
To quantitatively evaluate the effectiveness of LPO in finding missed optimization opportunities and to objectively compare it with the baselines, we conducted a controlled experiment to assess how effectively each tool identifies previously reported missed optimizations. For the benchmark suite, we collected 25 missed peephole optimizations from issues in the official LLVM GitHub repository tagged with both missed-optimization and llvm:instcombine labels. To avoid potential data leakage (Zhou et al., 2025; López et al., 2025), we only included issues created after the end of August 2024, which is the latest knowledge cut-off date among all LLMs listed in Table 1 except for Gemini2.5, the model excluded from this research question. For each benchmark, we collected the example suboptimal LLVM IR function provided in the issue and examined whether LPO, LPO-, Souper and Minotaur could detect the missed optimization within the function. To mitigate the effects of the inherent non-determinism of LLMs, we repeated each experiment five times for both LPO and LPO-. The results of the experiment are shown in Table 2.
Effectiveness of LPO with Different LLMs.  The effectiveness of LPO heavily depends on the code understanding and reasoning capabilities of the employed LLM, and LLMs with different capabilities can therefore lead to varying levels of effectiveness. As shown in Table 2, the performance of LPO varies significantly across different LLMs. With Gemma3, the smallest and weakest of the selected models, LPO can detect at most 3 out of 25 missed optimizations, and on average only 1.2 per round. With other non-reasoning but more capable models, the performance of LPO improves substantially: using Llama3.3, Gemini2.0, and GPT-4.1, LPO can identify up to 7, 11 and 12 missed optimizations, and on average 7.0, 7.4, and 7.4, respectively. Employing reasoning models further enhance the performance of LPO. With Gemini2.0T and o4-mini, LPO can detect up to 21 and 18 missed optimizations, and on average 15.6 and 14.2, respectively.
@@
LPO vs. LPO-.  The iterative prompting strategy improves LPO’s effectiveness in identifying missed optimizations regardless of the model employed. Without iterative prompting, LPO- failed to identify 1 to 7 optimization that LPO can detect across different LLMs. Moreover, the average number of optimizations detected by LPO- is consistently lower than that of LPO, with the gaps ranging from 0.4 to 5.2.
[…]
Issue ID Status $_{\texttt{Default}} Souper$$_{\texttt{Enum}}$ Minotaur
@@
154238 Fixed
154242 Confirmed
154246 Confirmed
154258 Unconfirmed
157315 Fixed
157370 Fixed
Issue ID Status $_{\texttt{Default}} Souper$$_{\texttt{Enum}}$ Minotaur
157371 Fixed
157372 Duplicate
157486 Confirmed
157524 Fixed
163084 Confirmed
163093 Unconfirmed
163108 Fixed
163109 Confirmed
163110 Confirmed
163112 Confirmed
163115 Confirmed
166878 Confirmed
166885 Confirmed
166887 Unconfirmed
166890 Unconfirmed
166973 Fixed
167003 Confirmed
167014 Confirmed
167055 Confirmed
167059 Unconfirmed
167079 Unconfirmed
167090 Unconfirmed
167094 Duplicate
167096 Confirmed
167173 Confirmed
@@ 4.3. RQ2: Discovering New Missed Optimization
To demonstrate the practical value of LPO, we intermittently ran an experiment that employs LPO to search for new missed optimizations in LLVM. In this experiment, the corpus used by LPO is from an LLVM IR dataset named LLVM Opt Benchmark (Zheng, 2023). This dataset contains optimized IR files from 240 real world projects written in C, C++, or Rust. Since the entire benchmark is too massive to serve as our corpus, we built our corpus by selecting five popular projects for each programming language, i.e., cpython, ffmpeg, linux, openssl, redis, node, protobuf, opencv, z3, pingora, ripgrep, typst, uv, zed. From this corpus, LPO extracted over 800,000 unique instruction sequences, eliminating around 8.7 million duplicates.
Table 3 lists all missed optimizations found by LPO and reported to LLVM. Within eleven months of development and intermittent testing, LPO has found 62 potential missed peephole optimizations, with 28 confirmed and 13 already fixed in LLVM. It also detected a bug in Alive2.222https://github.com/AliveToolkit/alive2/issues/1229 Among all the reported missed optimizations, we marked 3 of them as wontfix (i.e., Issue 130954, Issue 132628, and Issue 167199) according to the feedback from the maintainers. The reasons are: (1) the optimization is handled by the backend, (2) the optimization would prevent other valuable optimizations, and (3) the optimization is application-specific and is not general enough.
@@
Case Study 1 (Figures 4(a) and 4(d)).  The src function loads two consecutive 16-bit values from a given address %0 and combines them into a single 32-bit value using zero extension (zext), left shift (shl), and an or instruction. This function can be optimized to directly loading a 32-bit value from the given address. LLVM originally missed this optimization because it did not support merging consecutive loads with different sizes. Specifically, the original unoptimized function loads three consecutive values—two 8-bit values and one 16-bit value. LLVM successfully merges the two 8-bit loads but fails to further merge them with the subsequent 16-bit load. Souper and Minotaur cannot identify this case, as they cannot synthesize load and getelementptr instructions.
Case Study 2 (Figures 4(b) and 4(e)).  In the src function, the input %0 is first clamped to be at least 1 by llvm.umax.i8, then doubled with a 1-bit left shift, and finally clamped again to be at least 16. In the tgt function, the initial clamp is removed. This simplification is valid because the second clamp to 16 subsumes the effect of the first clamp, which makes it redundant. Souper cannot identify this missed optimization because it does not support the intrinsic family llvm.umax.*. Although Minotaur supports synthesizing this operation, it fails to detect the missed optimization.
Case Study 3 (Figures 4(c) and 4(f)).  In the src function, the input %0 is first checked for being an ordinary (non-NaN) value; if it is NaN, it is replaced with 0.0, otherwise, %0 is kept. The result is then compared against 1.0. In the tgt function, the check and replacement are removed, and %0 is directly compared with 1.0. This simplification is valid because the instruction fcmp oeq (ordered equal) already returns false if %0 is NaN, so explicitly substituting NaNs with 0.0 is redundant. Souper cannot identify this missed optimization because it does not support floating-point instructions, and Minotaur crashes on this IR function.
@@
double %0, double 0.000000e+00
%4 = fcmp oeq double %3, 1.000000e+00
@@
Figure 4. Three examples of confirmed missed optimizations found by LPO that Souper and Minotaur fail to detect.